# Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations

### v1

@@ Introduction
Related Work
Role‑Playing Agents. Early persona‑grounded dialogue work such as Persona‑Chat demonstrated that conditioning dialogue models on explicit character descriptions improves conversational coherence and engagement (Zhang et al. 2018). With the advent of LLMs, researchers have pursued finer‑grained persona consistency. Ji et al. (2025) propose a Persona‑aware Contrastive Learning framework that automatically aligns generations with the target character. CharacterGPT (Park, Park, and Lim 2025) proposes a framework that implements four fictional characters through eight attributes. DITTO (Lu et al. 2024) leverages character information to prompt instruction-following LLMs to to engage in role-playing dialogues, treating them as a form of reading comprehension. CharacterGLM (Zhou et al. 2024) introduces an open LLM family that aligns social characters with social traits
Complementary benchmarks soon followed, along with various evaluation methods. RPEval (Boudouri et al. 2025) evaluates emotional understanding, decision‑making, moral alignment, and in‑character fidelity in single‑turn settings, while TimeChara (Ahn et al. 2024) probes point‑in‑time role‑playing and exposes temporal hallucinations in state‑of‑the‑art LLMs. InCharacter (Wang et al. 2024b) adopts an interview-driven methodology to precisely assess RPA personalities by eliciting their underlying mindsets and behaviors. Wang et al. (2025) provides a dynamic, multi-agent virtual environment specifically designed to draw out nuanced, human-like behaviors from large language models during role-playing assessments.
However, current RPAs still truncate persona text, hallucinate beyond retrieved knowledge, and lack a benchmark for incremental updates—gaps our framework fills. In this paper, we propose a novel RAG-based role-playing framework that facilitates the addition and deletion of information.
Retrieval‑Augmented Generation (RAG). RAG couples non-parametric memory with an LLM to mitigate hallucination and stale knowledge (Lewis et al. 2020). Fixed‑length chunking can either omit critical context or waste prompt budget. Recent analyses quantify this trade‑off. Bhat et al. (2025) systematically evaluate fixed‑size chunking across multiple datasets and embedding models, showing that optimal sizes vary widely with domain and answer locality (Bhat et al. 2025). To move beyond one‑size‑fits‑all heuristics, Zhong et al. (2025) introduce Mix‑of‑Granularity, a router that dynamically selects chunk granularity per query and boosts medical QA accuracy (Zhong et al. 2025). Evaluation toolkits are likewise evolving. RAGEval generates scenario‑specific documents, QA pairs, and proposes Completeness, Hallucination, and Irrelevance metrics for rigorous RAG assessment (Zhu et al. 2024). Nevertheless, no existing benchmark explicitly targets role‑playing under RAG, and prior work still assumes personas are short, knowledge‑dense snippets. We mitigate this discrepancy with a novel text splitter that tailors chunk lengths and hierarchical context to each character. We also introduce CharacterRAG, the first dataset designed for building and evaluating RAG‑based role‑playing agents across 15 fictional personas.
Figure 2: An overview of our framework. The Adaptive Context-aware Text Splitter (ACTS) divides a persona into meaningful chunks, selecting the optimal chunk length and overlap for each character. Each chunk is further enhanced with hierarchical context, aiming to minimize information loss about the character. To answer questions beyond the character’s pre-existing knowledge, the Guided Selection (GS) module identifies chunks related to the relevant knowledge, while the Attribute Extractor (AE) extracts general attributes of the character from these chunks. Finally, our framework generates vivid responses by treating the chunks obtained from the RAG as knowledge and the outputs from GS and AE as the character’s attributes.
Task Formulation
Given user query $u$, RAG-based RPAs can be formulated as:
$\mathcal{R}=f(u,\mathcal{D}_{p})$ (1)
, where $\mathcal{D}_{p}$ is a character’s persona, and $f$ is a RPA. A text splitter $g$ divides the persona into $n$ chunks as follows:
$g(\mathcal{D}_{p})=\{c_{1},c_{2},...,c_{n}\}$ (2)
Each chunk $c_{i}$ contains the character’s attributes such as relationships and backstory. The objective of $f$ is to vividly embody the character and generate response $\mathcal{R}$ to $u$ while maintaining persona consistency. However, previous RAG methods (Guu et al. 2020; Guo et al. 2024; Yang et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025) truncate each character’s persona paragraph to a fixed length, regardless of the varying lengths across characters, which results in hallucinations or responses with lower persona consistency. In addition, if the persona does not contain knowledge relevant to the given query, RAG-based role-playing methods are prone to hallucination, making it challenging to generate accurate responses.
Figure 3: Statistics of CharacterRAG Dataset. CharacterRAG consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
Method
As shown in Figure 2, our framework consists of (i) Adaptive Context-aware Text Splitter (ACTS), (ii) Guided Selection (GS), and (iii) Attribute Extractor (AE), in order to build a realistic role-playing agent based on RAG.
Adaptive Context-aware Text Splitter
Given the character’s persona $\mathcal{D}_{p}$, ACTS aims to preserve the context between chunks, as well as the information about corresponding subsections within the persona for each chunk, that is, the hierarchical context. To this end, ACTS first finds the maximum length of the paragraphs that constitute the persona as follows:
$l_{\text{max}}=\varphi(p_{1},p_{2},...,p_{l})$ (3)
, where $\varphi$ denotes a length-calculating function. Then, ACTS sets the overlap length of the text splitter to half of $l_{\text{max}}$ (i.e., $l_{\text{o}}=l_{\text{max}}/2$). Note that the reason for setting the overlap length sufficiently large is to minimize information loss, as the context between pieces of information contained in each chunk is indispensable in RAG-based role-playing. Finally, $\mathcal{D}_{p}$ is segmented using $l_{\text{max}}$ and $l_{\text{o}}$ and hierarchical context $\mathcal{H}_{i}$ is appended to each chunk in order to preserve information such as descriptions of the corresponding characters and situational context at each point in the narrative:
$\text{ACTS}(\mathcal{D}_{p},\mathcal{H},l_{\text{max}},l_{\text{o}})=\{\hat{c}_{1},\hat{c}_{2},...,\hat{c}_{m}\}$ (4)
$\hat{c}_{i}=c_{i}+\mathcal{H}_{i}$ (5)
Guided Selection
@@
Input: User query $u$; Set of knowledge chunks $\mathcal{C}$; Maximum number of search iterations $N$; Slot size $K$
@@
2: Sort chunks $\mathcal{C}$ in descending order according to similarity to $u$, obtaining $\mathcal{C}_{\text{sorted}}$
@@
[…]

### v3

@@ 5.1 Setup
5.2 Evaluation Protocols
Tasks. As we follow the similar experimental protocol in prposed by previous studies (Wang et al., 2024b; Park et al., 2025), we also exploit 60 MBTI questions and 120 BFI (Barrick and Mount, 1991) questions to investigate whether each character can appropriately respond to questions for which they do not have explicit prior knowledge (Figure 4). Following the prior work, since it is not possible to construct QA pairs for questions outside the scope of a character’s knowledge, we instead conduct interview-based assessments (Wang et al., 2024b) for each character and compare the results to psychological test outcomes for the character, as determined by thousands of actual participants’ votes [^3] https://www.personality-database.com/(Sang et al., 2022; Wang et al., 2024b). We also use 450 QA pairs from the CharacterRAG dataset to verify whether the RPA sufficiently leverages each character’s knowledge.
Metrics. We design three LLM-based metrics, similar to those in prior studies (Wang et al., 2025a, 2024b), to comprehensively evaluate the role-playing capabilities of RAG-based RPAs.: (i) ACC measures whether the character’s response contains the correct answer or not. (ii) ACCL is a score assigned by the LLM, ranging from 1 to 10, that evaluates how well the character’s response reflects the correct answer. (iii) Hallucination Score (HS) evaluates the degree of hallucination in the model’s response given a query, the relevant chunks, and the ground-truth answer, on a scale from 1 to 10. Specifically, HS is assigned close to 1 when the response faithfully reflects only the facts contained in the chunks or answer without distortion or addition, indicating minimal hallucination.
Table 1: Distribution of Similarity Scores. We analyze the distributions of similarity scores to examine whether splitting the text into optimally sized chunks for each character’s persona and incorporating hierarchical context is effective.
[…]
$\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$) $\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$) $\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$)
RCTS 6.4325 0.1026 8.3306 0.1557 12.3136 0.0071
MHTS 6.4262 0.1038 8.3410 0.1552 12.3063 0.0071
SC 5.3405 0.1625 8.1691 0.1783 - -
ATS 6.7007 0.0884 8.4718 0.1281 12.2336 0.0070
ACTS (Ours) 6.8575 0.0784 8.6226 0.1179 12.3240 0.0063
Table 2: Human Evaluation. $S$ represents the results of all human evaluators, $\mu$ is $\mathbb{E}(\mathbb{E}(S))$, $\sigma$ is $\mathbb{E}(\sigma(S))$, and Mdn is $\mathbb{E}(Mdn(S)).$
$\mu$ ($\uparrow$) $\sigma$ ($\downarrow$) Mdn ($\uparrow$) Cronbach’s alpha ($\uparrow$)
BFI 3.970 0.962 4.217 0.825
MBTI 3.902 0.915 4.000 0.810
Table 3: Predicted MBTI Types per Character. Experiments are conducted using GPT-4.1 setting.
Naive RAG CRAG RAPTOR Adaptive RAG LightRAG AMADEUS (Ours) GT
Anya Forger ISFP (-2) INTJ (-3) INFJ (-2) INFJ (-2) INFJ (-2) ENFP (0) ENFP
Chika Fujiwara ENFP (0) ENFP (0) ENFJ (-1) ENFJ (-1) INTP (-2) ENFP (0) ENFP
Edward Elric INTP (-1) ISTJ (-3) INTP (-1) ENFP (-1) INTP (-1) INFP (-2) ENTP
Frieren INFP (-1) INFP (-1) INFP (-1) INFP (-1) INTP (0) INTP (0) INTP
Hitori Gotoh ISFP (-1) ISTJ (-2) INFP (0) INFP (0) ENFP (-1) INFP (0) INFP
Light Yagami INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) ENTJ
Māo Māo ISTJ (-2) INTJ (-1) ISTJ (-2) ENTJ (-2) INTJ (-1) ISTP (-1) INTP
Megumin ISFP (-1) INFP (0) ENFP (-1) ENFP (-1) INFP (0) ISFP (-1) INFP
Mikoto Misaka ENFP (-2) ISFP (-4) ENFJ (0) ENFJ (0) ENTJ (0) INFJ (-2) ENTJ
Nina Iseri INFP (-1) ISFP (0) ENFJ (-1) INFJ (-2) ENFJ (-3) INFP (-1) ISFP
Saitama ISFP (-1) ISTP (0) ISFJ (-2) INFJ (-3) INTP (-1) ISTP (0) ISTP
Son Goku ESFP (0) ENFJ (-2) ENFJ (-2) ENFP (-1) INTP (-3) ESFP (0) ESFP
Tanjiro Kamado ENFP (-1) ENFJ (0) ENFJ (0) ENFJ (0) INTP (-3) ENFJ (0) ENFJ
Tobio Kageyama ENFJ (-3) ENTJ(-2) ISTJ (0) ISTJ (0) INFJ (-1) ISTJ (0) ISTJ
Yui Hirasawa ISTJ (-4) ENFP (0) ENFP (0) ENFP (0) INTP (-2) ESFP (-1) ENFP
$\sum|d|$ ($\downarrow$) 21 19 14 15 21 9 -
Accuracy ($\uparrow$) 65.00% 68.33% 76.67% 75.00% 65.00% 85.00% -
Avg F1-Score ($\uparrow$) 0.6146 0.6448 0.6925 0.6832 0.5344 0.8244 -
Table 4: Predicted Big 5 SLOAN Types per Character. Experiments are conducted using GPT-4.1 setting.
Naive RAG CRAG LightRAG AMADEUS (Ours) GT
Anya Forger SLOAI (-1) SLOAI (-2) RCUEN (-3) SLUEI (-2) SCUAI
Chika Fujiwara SCOAI (-1) SCUAI (0) RLUEN (-4) SCOAI (-1) SCUAI
Edward Elric SCOAI (-3) SLOEI (-1) RCUAN (-4) SLOEI (-1) SLUEI
Frieren RCOAI (-2) RCUAI (-1) SLUEN (-3) RCUAI (-1) RCUEI
Hitori Gotoh RLUAI (0) RLUAI (0) RCUEN (-3) RLUAI (0) RLUAI
Light Yagami SCOEI (-1) SCOEI (-1) RCUAN (-3) SCOEI (-1) RCOEI
Māo Māo RCOEI (0) RCOAN (-2) SLUAN (-5) RCOEI (0) RCOEI
Megumin SCOAI (-3) SCUAI (-2) RLUEN (-2) SLOEI (-1) SLUEI
Mikoto Misaka SLOAI (-3) SLOAI (-3) RCUEN (-2) SLOAI (-3) RCOEI
Nina Iseri SLOAI (-3) RLUAI (-1) RCUEN (-2) SLUEI (-1) RLUEI
Saitama RCUAN (0) RCOAN (-1) SCOAI (-3) RCUAN (0) RCUAN
Son Goku SCOAI (-2) SCUAI (-1) RLUEI (-4) SCOAI (-2) SCUAN
Tanjiro Kamado SLOAI (-1) SCOAI (0) RCUAN (-3) SCOAI (0) SCOAI
Tobio Kageyama RCOEN (-1) SLOAN (-2) RCUAI (-4) RCOEN (-1) RLOEN
Yui Hirasawa SCUAI (0) SLUAI (-1) RCOEN (-4) SCUAI (0) SCUAI
$\sum|d|$ ($\downarrow$) 21 18 49 14 -
Accuracy ($\uparrow$) 72.00% 76.00% 34.67% 81.33% -
Avg F1-Score ($\uparrow$) 0.6785 0.7313 0.2774 0.7986 -
Table 5: Ablation Study. We analyze the performance gain obtained from incrementally adding each substage.
Method $\sum|d|$ ($\downarrow$) Accuracy ($\uparrow$) Avg F1-Score ($\uparrow$)
Naive RAG 21 65.00% 0.6146
Naive RAG + ACTS 19 68.33% 0.6524
Naive RAG + ACTS + GS 15 75.00% 0.7426
Naive RAG + ACTS + GS + AE (Ours) 9 85.00% 0.8394
Table 6: Role-playing capabilities on CharacterRAG. Higher values of ACC (%) and ACCL (1-10) correspond to better performance, whereas lower values of HS (1-10) are preferable.
RAG Method GPT-4.1 Gemma3-27B Qwen3-32B
ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$) ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$) ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$)
w/o RAG 49.56% 6.79 - 27.56% 5.33 - 18.89% 4.35 -
Naive RAG 91.33% 9.23 3.13 86.44% 8.85 3.27 78.44% 8.49 5.05
LightRAG 48.00% 6.06 - 69.56% 8.17 - 68.67% 8.20 -
CRAG 70.00% 8.26 3.21 57.78% 7.57 4.09 28.67% 5.24 8.68
AMADEUS (Ours) 92.67% 9.26 2.89 88.00% 8.92 3.26 78.89% 8.63 4.66
5.3 Experimental Results
Adaptive Persona Segmentation and Hierarchical Contextualization Are Highly Effective. In Table 1, we provide each character with 30 questions, resulting in a total of 450 questions, related to their respective knowledge from the CharacterRAG dataset, and measure the similarity between each question and the chunks retrieved by the RAG model under the three different embedding settings. Results demonstrate that compared to RecursiveCharacterTextSplitter (RCTS) (LangChain, 2024b), MarkdownHeaderTextSplitter (MHTS) (LangChain, 2024a), and SemanticChunker (SC) (LangChain, 2024c), adaptive persona segmentation, which we call Adaptive Text Splitter (ATS), segments text with an optimal persona length and overlap for each character and without hierarchical context, achieves a higher average score and lower variance. This indicates that each chunk generated using adaptive persona segmentation contains richer semantic information for the same query. Building on this, ACTS, which considers hierarchical context in addition to ATS, consistently achieves better performance across all three embedding settings. These results show that optimal chunk length, appropriate overlap, and consideration of hierarchical context all play essential roles in effective text chunking.
Extracting a Character’s Attributes from Selected Chunks Is Reliable. We investigate the reasonableness of inferring character attributes with AE from chunks extracted via GS by conducting human evaluation using a 5-point Likert scale. To this end, we invite 14 human evaluators and each evaluator is asked to score 60 randomly selected samples. Each sample consists of pairs of chunks selected from GS and attributes extracted through AE, for 30 BFI questions and 30 MBTI questions. Table 2 shows that the means $\mu$ is close to 4, with small standard deviations $\sigma$. It demonstrates that the outputs of GS and AE are reliable and trustworthy, even from a human evaluative perspective. We measure Cronbach’s alpha (Cronbach, 1951) to evaluate internal consistency among the human evaluators. Note that the Cronbach’s alpha values are 0.825 and 0.810, both exceeding the commonly accepted threshold of 0.7 for acceptable reliability.
Graph-Based and Web Search-Based RAG Are Unsuitable for RAG-Based RPAs. One of the major challenges in retrieval-based role-playing is that, when a RPA receives questions involving knowledge outside a character’s persona, it tends to either overuse irrelevant chunks (Figure 6) or generate uninformative responses (Guo et al., 2024; Shukla et al., 2025; Wang et al., 2025b). To investigate whether RAG-based RPAs can handle this problem, we conduct extensive experiments in which we ask 15 characters 60 MBTI questions and 120 BFI questions each, and and evaluate their ability to accurately infer the characters’ personality types. Table 3 and Table 4 shows predicted MBTI types and Big 5 SLOAN types per character. The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified. $\sum|d|$ is a measure obtained by summing these values; lower values are preferable. Our framework maintains persona consistency even when answering questions that are not explicitly specified in each character’s persona in both MBTI and BFI settings. Note that the performance gap of CRAG is significant between the two settings. We assume that questions requiring analogical reasoning are difficult to solve even with web search and that the search results may contain non-negligible noise. On the other hand, LightRAG exhibits the lowest performance, which shows that graph-based RAG methods are not well suited for RPA applications due to the high cost of graph construction, the difficulty in adding or removing new knowledge, and challenges in maintaining persona consistency. While we did not perform a direct comparison, we observed that GraphRAG (Shukla et al., 2025) suffers from similar problems.
Impact of Each Module in Our Framework. Table 5 demonstrates the results of ablation study, evaluated on 60 MBTI questions. We find that ACTS, GS, and AE are all highly effective at handling questions that go beyond a character’s knowledge. Note that AE achieves the largest performance improvement, highlighting that extracting and incorporating character attributes as input is far more effective than simply providing chunks related to the user query.
CharacterRAG Dataset Serves as a Valuable Resource for the Construction and Evaluation of RAG-Based RPAs. To investigate the factors influencing the performance of role-playing, we conduct a comprehensive interview-based assessments on the generalization capabilities of models with various LLMs and RAG techniques. Table 6 and Table 8 present how the ability to accurately answer questions related to the character’s knowledge, which is a core aspect of role-playing, varies across the applied methodologies. We first examine to what extent each LLM possesses background knowledge about the 15 characters in a setting without RAG, and results show that none of the three LLMs are capable of effective role-playing without access to external knowledge. Moreover, we observe that LightRAG, a graph-based RAG, is ill-suited for the storage and retrieval of character knowledge, as it often suffers from issues such as entity ambiguity and uninformative responses.
In a similar vein, CRAG exhibits challenges in maintaining role-playing fidelity, which can be attributed to the tendency of web search-based RAG methods to utilize retrieved content that may undermine the consistency of a character’s persona. Indeed, despite leveraging web information, CRAG is able to correctly answer only 6 out of the 30 CharacterRAG questions pertaining to Nina Iseri. In addition, to analyze how a thinking mode of LLMs influences their role-playing capabilities, we employ Qwen 3-32B. Results demonstrate that the thinking mode fails to yield any substantial positive effect on enhancing role-playing performance. Our framework achieves the best performance across all three LLMs. We also find that the Hallucination Score (HS) is the lowest in CharacterRAG setting.
6 Conclusion
Ethical Considerations
This study follows ethical guidelines to ensure the integrity and fairness of all experiments.
Data and Privacy. The Namuwiki portion of our data was gathered within the bounds of the platform’s licensing/usage rules. We verified that the collection contained no personally identifiable information, and we restricted its use to academic research. For personas and prompting, we relied on either publicly released content, anonymized artifacts, or content created specifically for this study with ethical considerations in mind.
Experimental Design. Our experiments were designed in accordance with rigorous ethical guidelines, with particular emphasis on safeguarding participant data and maintaining privacy. To mitigate potential bias, we recruited evaluators with diverse backgrounds and carried out the human evaluation process using procedures intended to be equitable and transparent. All participants received clear information regarding the study’s objectives and procedures, and they were allowed to discontinue their participation at any point without negative consequences. Through compliance with these practices, we seek to advance AI research in a way that is both scientifically meaningful and ethically grounded, prioritizing privacy, intellectual property rights, and the welfare of everyone involved.
Appendix A Chunk Duplication Frequencies
Table 7: Work Genres in Which Each Character Appears. We list the work genres associated with the 15 characters included in the CharacterRAG dataset.
Character Genre
Anya Forger Action, Comedy, Drama, Slice of Life, Spy
Chika Fujiwara Gag, Romantic Comedy
Edward Elric Adventure, Fantasy
Frieren Fantasy, Adventure
Hitori Gotoh Band, Comedy, Drama, Music, Slice of Life
Light Yagami Crime, Dark Fantasy, Noir, Thriller
Megumin Comedy, Fantasy, Isekai
Mikoto Misaka Cyberpunk, SF, Slice of Life
Nina Iseri Band, Drama
Saitama Action, Comedy, Hero
Son Goku Battle, Comedy, Fantasy
Tanjiro Kamado Action, Adventure, Historical Fantasy
Tobio Kageyama Sports, Volleyball
Yui Hirasawa Band, School, Slice of Life
@@ Appendix A Chunk Duplication Frequencies
Input: User query $u$; Set of knowledge chunks $\mathcal{C}$; Maximum number of search iterations $N$; Slot size $M$
@@
2: Sort chunks $\mathcal{C}$ in descending order according to semantic similarity to $u$, obtaining $\mathcal{C}_{\text{sorted}}$
@@
15:  $S\leftarrow$ Top-$K+1$ chunks from $\mathcal{C}_{\text{sorted}}$ (highest semantic similarity to $u$)